OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote of y = 2/9

OpenStudy (jennychan12):

y = 1/9x-2

OpenStudy (anonymous):

That doesnt help :/

OpenStudy (anonymous):

just choose a rational function where the degree for the numerator equals the degree of the denominator and have the leading coefficients be 2 and 9, respectively.

OpenStudy (anonymous):

still incorrect.

OpenStudy (anonymous):

if you follow my instructions, it will be correct....

OpenStudy (anonymous):

Well from my understanding is that when you input the values that the asymptotic is you want to make the function undefined. One way that happens is when you have something divided by a zero. \[\frac{1}{x+?}\] x is for vertical asymptote and y is for horizontal asymptote \[\frac{1}{y+?}\] so we use the y and we set it equal to the other unknown. \[\frac{1}{y+?}=x\] The question mark is the negative value for where the asymptote occurs so since it occurs in 2/9 the question mark will equal -2/9 \[\frac{1}{y-\frac{2}{9}}=x\]

OpenStudy (anonymous):

Usually we have it set to y so you have to mess around with it until we get it in the form we want. That's why I like verticals asysmptote because those are already set to y.

OpenStudy (anonymous):

If you want to see if there is a horizontal asymptote then all you have to do is graph it and you can use a calculator or wolfram alpha .com